OpenStudy (anonymous):

Solve for x. 2 • 3 – 2 • 11 = 2(x – 11)

OpenStudy (hlambach):

Simplify 64 = 2x -11 Add 11 to each side 75 = 2x Divide each side by 2 37.5 -x

OpenStudy (anonymous):

=6-22=2x-22 so 2x=6 after adding 22 on both sides so x=3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome

OpenStudy (hlambach):

yep

OpenStudy (hlambach):

whoops i did it wrong the other has it right sry!

OpenStudy (hlambach):

haha

OpenStudy (anonymous):

its ok i got it trhx for trying ;)

OpenStudy (hlambach):

:)

OpenStudy (hlambach):

forgot to multiply 11... oh well :)